OpenStudy (rawfeeds-):

The coordinates below represent two linear equations. How many solutions does this system of equations have? Line 1 x y –3 5 0 4 Line 2 x y 0 4 6 2 A. 0 B. exactly 1 C. exactly 2 D. infinitely many

Arcadiouse (arcadiouse):

Show a picture of the graph please.

OpenStudy (rawfeeds-):

there is no graph

OpenStudy (rawfeeds-):

@Arcadiouse

Arcadiouse (arcadiouse):

Could you please write both the equations like they should be?

OpenStudy (rawfeeds-):

i need help

OpenStudy (kropot72):

One way to solve this is to find the gradient of each line. If the gradients are different, there is one solution. If the gradients are the same, there is no solution.

OpenStudy (wolf1728):

Line 1 slope = 5-4 / -3 -0 = 1/-3 Line 2 slope = 4-2 / 0 -6 = 2 / -6 = 1/-3 Their slopes are equal answer is D infinitely many.

OpenStudy (kropot72):

The slopes are equal. Therefore the lines never cross, so no solution is possible.

OpenStudy (wolf1728):

The gradients are the same and there are infinitely many solutions: X=-3 Y=3 X=0 Y=4 X=3 Y=5 X=6 Y=6 as well as MANY other solutions.

OpenStudy (kropot72):

@wolf If you plot the given co-ordinates, you will find that the "two linear equations" are actually one equation. Notice that one pair of co-ordinates for the "two linear equations" is exactly the same in each case. Therefore there would be an infinite number of solutions for a single linear equation. The question is misleading in referring to "this system of equations". Of course there is also the possibility that there was an error in the details of the co-ordinates in the question.